40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Mar 10 16:29:32 2016, gautam, Update, endtable upgrade, Inventory check Inventory_check.pdfDoubler_comparison.pdf
    Reply  Thu Mar 10 16:32:45 2016, rana, Update, endtable upgrade, Inventory check 
       Reply  Mon Mar 14 22:42:23 2016, gautam, Update, endtable upgrade, Inventory check 
       Reply  Mon Mar 28 10:43:18 2016, Steve, Update, endtable upgrade, ETMX 4'x2' optical table pictures ETMX4x2Layout.jpgETMX_4x2.JPGETMXopt_4x2.JPG
          Reply  Wed Mar 30 17:38:52 2016, gautam, Update, endtable upgrade, X end table proposed layout  ETMX_3x2.JPGlayout_details_20160328.pdfETMX_proposed_layout.pdf
             Reply  Thu Mar 31 09:38:41 2016, Steve, Update, endtable upgrade, Y end 4x3 existing layout ETMY4x3layout.jpg
                Reply  Thu Mar 31 19:49:31 2016, gautam, Update, endtable upgrade, proposed layout v2 ETMX_proposed_layout_v2.pdf
                   Reply  Mon Apr 4 10:59:12 2016, gautam, Update, endtable upgrade, proposed layout v3 ETMX_proposed_layout_v3.pdfIR_modematch.pdfGreen_modematch.pdfXendALaMode.zipETMX_proposed_layout_dimensions.dwg
                      Reply  Mon Apr 4 15:04:14 2016, gautam, Update, endtable upgrade, COMPONENT REMOVAL 
                         Reply  Tue Apr 5 11:42:17 2016, gaericqutam, Update, endtable upgrade, TABLE REMOVAL 
                            Reply  Wed Apr 6 17:52:21 2016, gautam, Update, endtable upgrade, First contact cleaning commenced 
                               Reply  Thu Apr 7 12:51:24 2016, gautam, Update, endtable upgrade, Beam height differences 
                                  Reply  Thu Apr 7 14:30:58 2016, Steve, Update, endtable upgrade, ETMX table height ETMXbrs.pngETMX.jpgIMG_0618.JPGETMX.jpgETMXwc.jpg
                                     Reply  Fri Apr 8 09:13:20 2016, Steve, Update, endtable upgrade, ETMX-T beam height ETMXalignedT.pngETMXF.jpg
Message ID: 12056     Entry time: Wed Mar 30 17:38:52 2016     In reply to: 12051     Reply to this: 12057
Author: gautam 
Type: Update 
Category: endtable upgrade 
Subject: X end table proposed layout  

Attachment 1: This is a photo of the current X end table optical layout with the beampaths of the various sub-systems overlaid. For the labels, see Attachment #2.

Attachment 2: This is a summary of all the optical components that are currently being used. I've noted some things we may want to change when we effect the swap. The important ones are:

  • Switch out all 1" and 2" optic mounts which are not of the Polaris type to the Polaris type. I have checked that we have sufficient numbers of these in hand.
  • Adjust the collimating lens of the fiber collimating telescope to get a better mode
  • Many of the labels are probably outdated, now would be a good time to update them
  • For the mode-matching of the AUX IR into the doubling crystal, a la mode suggests a better (i.e. less sensitive to lens position) solution is effected with L2 as a 100mm fl lens rather than 88.3mm. I did not change this during the laser swap in order to minimize the number of components changed. Since we are doing a wholesale change now, it may not be a bad idea to swap this out as well. I have checked that we have a suitable AR1064 coated lens.
  • Some optics probably need to be cleaned...
  • PZT mirror 2 has a new mount ready that is the "correct" height so we don't have to keep using makeshift stacked posts.
  • The plan as it stands is to use the green coloured mount for the IR faraday (IO-5-1064-HP).

Have I missed anything important?

Attachment #3: I've made a CAD drawing of the proposed new layout and have overlaid the beampath in an amateur way because I couldn't figure OptoCad out - I figure this will suffice for now. I have adopted elements from the current Y-end layout, but have used Anders' mode-matching solution (same lenses, same positions of optics) to make sure we have good Guoy phase separation between the two PZT steering mirrors. Some notes:

  • I've tried to palce the optics for the AUX IR into the doubler and subsequent steering of green into the arm cavity as per the mode matching solution. These should be pretty accurate, and the layout suggests we have some room to maneuver 
  • The Green REFL beampath is exaggerated but I think we have enough room to place Y16 appropriately and steer the reflected beam into the PDA36A
  • We need two more 1" 1064nm coated mirrors for the initial steering into the doubling oven, I have checked we have these in hand.
  • The IR pickoff into the fiber coupler may change somewhat once we change the mode and redo the mode-matching calculations. But again, I think we have sufficient room to implement a workable solution.
  • After accounting for the fact that the new endtable will be a little closer to the vacuum chamber, Y12 in the proposed layout will be ~10cm further away from ETMX than it is currently. But as discussed at the meeting today, the Rayleigh range of the green beam should be large enough here such that this shouldn't be a significant change.

Steve says the table is ready - so if we are happy with this layout, we can move forward...

Attachment 1: ETMX_3x2.JPG  2.020 MB  | Show | Hide all | Show all
Attachment 2: layout_details_20160328.pdf  28 kB  | Hide | Hide all | Show all
layout_details_20160328.pdf
Attachment 3: ETMX_proposed_layout.pdf  299 kB  | Hide | Hide all | Show all
ETMX_proposed_layout.pdf
ELOG V3.1.3-